# Elevated UV luminosity density at Cosmic Dawn explained by non-evolving, weakly-mass dependent star formation efficiency

### v1

@@ 3.2 The star formation efficiency of galaxies
The star formation efficiency (SFE) is a dimensionless quantity that links the star formation activity of a galaxy to the mass growth rate [^6] The SFE is sometimes defined relative to the baryonic accretion rate of halos. In this case the SFE is larger by a factor $1/f_{\rm bar,uni}$ ($\sim{}0.8$ dex) compared to the definition used in this work, where $f_{\rm bar,uni}\sim{}0.157$ is the universal baryon fraction. of its halo. In this work we consider both time-averaged SFRs as well as effective SFRs, derived from both the intrinsic and the dust-attenuated UV luminosity, to quantify star formation activity. Average halo growth rates are calculated based on the current halo mass, using the approximation by Behroozi & Silk (2015), both for simplicity and to allow for a better comparison with the literature. The goal of this section is to characterize the SFE as function of halo mass and redshift as predicted by FIREboxHR.
@@
where $\epsilon$ is a (mass dependent) efficiency factor and $t_{\rm ff}\approx{}0.06\,H^{-1}(z)$ is the halo free fall time, see Ferrara et al. (2023). We can motivate this equation on dimensional grounds and via the equilibrium ansatz (Bouché et al., 2010; Davé et al., 2012) balancing gas inflow into galaxies with star formation and outflows from galaxies. In equilibrium, $\dot{M}_{\rm gas,in}=(1-R+\eta{})\,{\rm SFR}$, where $\dot{M}_{\rm gas,in}$ is the gas inflow rate into galaxies, $R$ is the mass return fraction from evolving stellar populations in the instantaneous recycling approximation, and $\eta$ is the mass loading factor of galactic outflows. Provided the gas inflow rate into galaxies scales proportional to $M_{\rm halo}/t_{\rm ff}$, we recover eq. (4) with $\epsilon\propto{}1/(1-R+\eta{})\sim{}1/\eta$ for low mass galaxies. The alternative choice $\dot{M}_{\rm gas,in}\propto{}\dot{M}_{\rm halo}$ is discussed further below.
@@ 3.3 The stellar mass – halo mass relation
One interesting aspect is the scaling of the predicted stellar fraction with halo mass at the massive end. Provided $M_{\rm star,min}\ll{}M_{\rm star}$, the halo mass scaling of the SFE is closely related to the scaling of the SHMR. Specifically, for a power-law ${\rm SFE}\propto{}M_{\rm halo}^{\alpha}$ (with $\alpha>-1$), $M_{\rm star}\propto{}M_{\rm halo}^{1+\alpha}$ and the stellar fraction $f_{\rm star}=M_{\rm star}/M_{\rm halo}\propto{}M_{\rm halo}^{\alpha}$, i.e., the ‘integrated SFE’ $f_{\rm star}$ has the same scaling with halo mass as the instantaneous SFE. A broken power-law fit to the predicted $f_{\rm star}$ – halo mass relation returns a high mass slope of $0.33$, in line with the expectation from the halo mass scaling of the SFE. This slope is, however, lower than the high mass slope ($\sim{}0.50$) of the actual $f_{\rm star}$ – halo mass relation in FIREboxHR.
@@ 3.4 A theoretical model for the UV luminosity function and density at $z>6$
As discussed in Section 3.1, FIREboxHR predicts UV luminosity densities at $z>10$ in good agreement with observational data by JWST. In this section, we introduce a simple model that provides a theoretical explanation for this result. Our approach builds on the ideas of previous work (Trenti et al., 2010; Tacchella et al., 2013; Mason et al., 2015; Furlanetto et al., 2017; Tacchella et al., 2018; Harikane et al., 2022; Lovell et al., 2023; Ferrara et al., 2023) but differs in its use of a SFE – halo mass relation derived from a cosmological, hydrodynamical simulation. Given that the FIRE-2 physics model is not explicitly tuned to specific observables, this model is theoretical, and not empirical, in nature. This theoretical framework demonstrates that the gradual evolution of the UV luminosity density can be directly attributed to the baryonic processes, including stellar feedback and gas accretion, that take place in galaxies at high redshifts.
@@
It follows from eq. (9) that $\mathcal{L}$ is also a log-normally distributed random variable. In contrast, the rest-frame UV magnitude defined as
@@
where $p$ is the pdf of $\mathcal{M}$ with parameters given by (12). We can evaluate the integrals in eqs. (13) and (15) numerically with halo mass functions (HMFs) provided by the hmf module for Python (Murray et al., 2013; Murray, 2014).
The discussion so far has been quite general. The non-evolving SFE model makes the additional assumption, based on the results of the previous section, that the average SFE and its scatter in halos of a given mass do not vary with redshift at $z>6$.
Figure 6: Evolution of the ultraviolet (UV) luminosity density down to a limiting observed rest-frame magnitude of -17 as predicted by FIREboxHR (blue solid line and error bars) and by a theoretical model with a non-evolving star formation efficiency (SFE) – halo mass relation inferred from the simulation (red line and shaded areas). The red line shows the UV luminosity density in galaxies brighter than -17 as predicted by the theoretical model for the same halo mass distribution as in FIREboxHR, while dark and light shaded areas indicate how the scatter in the SFE affects the UV luminosity in the given volume at the 68 and 95 percent confidence level. Uncertainties (68% confidence level) of the FIREboxHR predictions are calculated via bootstrapping and account for both the scatter in SFE and variations in halo numbers. Long and short dashed lines reproduce the predictions of empirical models by Tacchella et al. (2018) and Harikane et al. (2022), and symbols with error bars show observational estimates (Bouwens et al., 2022; Donnan et al., 2023a; Donnan et al., 2023b; Harikane et al., 2023). The theoretical model reproduces the evolution of the UV luminosity density found by FIREboxHR within the uncertainties.
@@
The theoretical model introduced in the previous section allows us to study the evolution of the UV luminosity density and UV LF during reionization and late Cosmic Dawn without being limited by simulation volume. To integrate eqs. (13, 15), we use the halo number densities provided by hmf version 3.4.4 for a Planck 2018 cosmology (Aghanim et al., 2020) with $h=0.6766$, $\Omega_{0}=0.3111$, $\Omega_{\rm b}=0.04897$, and $\sigma_{8}=0.8102$, a CAMB transfer function with [^8] Setting $k_{\rm max}$ to a sufficiently large value is critical as hmf’s default setting, at least in the code version we used, overestimates (e.g., by about 0.5 dex for $10^{8}\,M_{\odot}$ halos at $z=10$) the number density of low mass halos at high redshift. $k_{\rm max}=150$, and the Tinker et al. (2008) halo mass function with the Behroozi et al. (2013) extension to high redshift. To approximately account for the impact of baryonic effects on the halo mass (e.g., Sawala et al. 2013; Velliscig et al. 2014; Cui et al. 2014; Schaller et al. 2015; Beltz-Mohrmann & Berlind 2021), we reduce the halo masses provided by hmf by 0.07 dex, i.e., $M_{\rm halo}\rightarrow{}M_{\rm halo}^{\prime}=10^{-0.07}M_{\rm halo}$ and $n(>M_{\rm halo})\rightarrow{}n^{\prime}(>M_{\rm halo}^{\prime})=n(>M_{\rm halo})$. For the fiducial set up, we adopt the non-evolving SFE – halo mass relation with a mass-dependent scatter and with parameters provided in Tables 1 and 2 for dust-attenuated UV luminosities and halo mass binning. Variations of this set-up will be discussed in the text when needed. We stress that some model predictions involve extrapolations of the non-evolving SFE – halo mass ansatz to lower halo masses and higher redshifts than reliably probed by FIREboxHR. Additionally, the predictions assume no fundamental changes in baryonic physics and star formation processes at high $z$, such as contributions from Population III stars (see Section 3.7).
Figure 7: Predictions of the theoretical model for the ultraviolet (UV) luminosity density at $z\sim{}6-20$. The predictions assume a non-evolving star formation efficiency – halo mass relation with mass-dependent scatter with parameters provided in Tables 1 and 2, a Tinker et al. (2008) halo mass function with the Behroozi extension to high redshift (Behroozi et al., 2013), and Planck 2018 cosmology (Aghanim et al. 2020). Different lines correspond to different limiting magnitudes (see legend). Symbols show observational results for $M_{\rm UV,lim}=-17$ (Bouwens et al., 2022; Donnan et al., 2023a; Donnan et al., 2023b; Harikane et al., 2023; Finkelstein et al., 2024; Donnan et al., 2024). The dotted line is the UV luminosity density needed for a 21 cm absorption signal in the ‘minimal coupling’ regime (Madau, 2018). The UV luminosity density predicted by the theoretical model matches current observations and shows a steeper evolution at the highest redshift.
Fig. 7 presents the UV luminosity density and UV LFs predicted by the theoretical model for $z\sim{}6-20$. The model accurately captures the observed evolution of the UV luminosity density over $z\sim{}6-13.5$. Notably, $\log{}\rho_{\rm UV}$ does not evolve linearly with redshift at $z\geq{}6$ (cf. Ceverino et al. 2024) showing instead a steeper decline in UV luminosity density towards higher redshifts. For instance, the UV luminosity density at $z=16$ (for a limiting magnitude of -17) is predicted to be about two orders of magnitude lower than at $z=11$, highlighting a challenge for observational studies targeting galaxies at those redshifts (e.g., Robertson et al. 2023a). Extending this model speculatively to $z=20$ suggests a further reduction in luminosity density by an additional two orders of magnitude. Recent tentative observational evidence may suggest a sudden steepening of the UV luminosity density evolution between $z\sim{}12.5$ and $14.5$ (Robertson et al., 2023a; Donnan et al., 2024). While a steeper evolution at higher $z$ is in line with our findings, we note, however, that our model predicts a gradual steepening over $z\sim{}6-15$, not a sudden transition between two separate evolutionary regimes.
@@
Figure 8: Predictions of the theoretical model for the ultraviolet luminosity functions (UV LFs) at $z\sim{}6-20$. All predictions assume a non-evolving SFE – halo mass relation, a Tinker et al. (2008) halo mass function with the Behroozi extension to high redshift (Behroozi et al., 2013), and Planck 2018 cosmology (Aghanim et al. 2020). Solid lines of various colors indicate the model predictions for different redshifts. Dot-dashed and double dot-dashed lines show observational estimates of the UV LF at $z\sim{}6-12$ (Bouwens et al., 2022; Harikane et al., 2022; Donnan et al., 2023a; Harikane et al., 2023). (Left) Predictions for the fiducial model with a mass dependent scatter, see Table 2. Dotted lines (dashed lines) show fits of the model UV LFs to a modified Schechter (a double power law) function (Bouwens et al., 2017; Harikane et al., 2022). While both fitting functions describe the UV LFs of the theoretical model reasonably well, a double power law provides a significantly better fit, especially at $z\leq{}14$. (Right) Similar to left panel but for a mass-independent scatter of the SFE – halo mass relation. Solid (dotted, dashed) lines correspond to a scatter of 0.4 (0.2, 0.6) dex. A larger scatter increases the number density of UV luminous galaxies. In addition to a rapid decline in overall normalization, the theoretical model predicts a steepening in the faint-end of the UV LFs towards higher redshift. Fitting parameters for a modified Schechter function and double power law are provided in appendix D.
@@
The UV LFs predicted by the theoretical model are in approximate agreement with observations, see Fig. 8. Deviations at the luminous end partly arise due to the lack of massive halos in FIREboxHR (given its modest box size) needed to constrain the high mass behavior of the SFE – halo mass relation. In addition, it is possible that FIREboxHR underestimates the scatter in more massive halos. A larger scatter would indeed increase the number density of UV luminous galaxies as shown by the figure (see also Sun et al. 2023c). For instance, a scatter of 0.6 dex (compared with the lower $\sim{}0.3$ dex found in FIREboxHR for $M_{\rm halo}\sim{}10^{10}-10^{11}\,M_{\odot}$ halos) would be sufficient to match the reported UV LF at $z=12$ by Harikane et al. (2023). However, applying the same scatter to $z=10$ would then lead to an overprediction of the observed UV LFs at this earlier redshift (Donnan et al., 2023a). A peculiar feature of the model with mass-dependent scatter is that it introduces a pronounced change in slope of the UV LF at moderately faint luminosities (e.g., near $-16.5$ at $z=6$) whereas the UV LFs of the fixed-scatter models lack this feature.
We fit the predicted UV LFs both with a modified Schechter function, see Section 4.3 in Bouwens et al. (2017), and a double power law (e.g., Harikane et al. 2022). The modified Schechter function includes a multiplicative term $10^{-0.4\delta{}(M_{\rm UV}+16)^{2}}$ to model a roll-over of the LF at the faint end. The fitting parameters are provided in appendix D. While not expected a priori, our model predictions are well fit by either functional form. In either case, and independently of the chosen scatter of the SFE – halo mass relation, fits to the UV LFs confirm the visually apparent decrease in the overall normalization and the steepening in the faint end slope with increasing redshift.
@@
Finally, the figure analyzes how scatter in the SFE – halo mass relation, and the corresponding variability in the UV magnitudes, affects the cosmic UV luminosity density and the UV LFs. Specifically, we vary the mass-dependent scatter by ±$\pm{}0.2$ dex finding that the cosmic UV luminosity density is largely insensitive to this scatter. This finding is in qualitative agreement with a recent study by Gelli et al. (2024) who concluded that scatter alone does not explain the high UV luminosity density at $z>12$. Scatter in the SFE – halo mass relation has a significant impact, however, on the bright end of the UV LFs. As discussed before, a higher UV variability boosts the number density of UV bright galaxies given their much lower number density compared to faint ones. However, unless the scatter is increased to a value much larger than found in FIREboxHR, the number density of galaxies which dominate the cosmic UV luminosity density is only weakly affected.
Our finding is in contrast to alternative models which aim to explain the comparably shallow evolution of the UV luminosity density at high $z$ with a large and redshift-dependent scatter of the SFE (e.g., Shen et al. 2023b). As we noted in Section 3.4, a parametrization of the SFE – halo mass relation based on the median SFE implies that an increase in the scatter also lead to an increase in the average SFE, thereby boosting the luminosities of all galaxies. A scatter that increases with redshift in this parametrization thus results in a redshift-dependence of the normalization of the average SFE – halo mass relation. In contrast, our theoretical model appears to match the observed evolution of the UV luminosity density without such an explicit redshift dependence.

### v2

@@ 3.2 The star formation efficiency of galaxies
The star formation efficiency (SFE) is a dimensionless quantity that links the star formation activity of a galaxy to the mass growth rate [^6] The SFE is sometimes defined relative to the baryonic accretion rate of halos. In this case, the SFE is larger by a factor $1/f_{\rm bar,uni}$ ($\sim{}0.8$ dex) compared to the definition used in this work, where $f_{\rm bar,uni}\sim{}0.157$ is the universal baryon fraction. of its halo. In this work we consider both time-averaged SFRs as well as effective SFRs, derived from both the intrinsic and the dust-attenuated UV luminosity, to quantify star formation activity. Average halo growth rates are calculated based on the current halo mass, using the approximation by Behroozi & Silk (2015), both for simplicity and to allow for a better comparison with the literature. The goal of this section is to characterize the SFE as function of halo mass and redshift as predicted by FIREboxHR.
@@
where $\epsilon$ is a (mass dependent) efficiency factor and $t_{\rm ff}\approx{}0.06\,H^{-1}(z)$ is the halo free fall time (Ferrara et al., 2023; Ferrara, 2024). We can motivate this equation on dimensional grounds and via the equilibrium ansatz (Bouché et al., 2010; Davé et al., 2012) balancing gas inflow into galaxies with star formation and outflows from galaxies. In equilibrium, $\dot{M}_{\rm gas,in}=(1-R+\eta{})\,{\rm SFR}$, where $\dot{M}_{\rm gas,in}$ is the gas inflow rate into galaxies, $R$ is the mass return fraction from evolving stellar populations in the instantaneous recycling approximation, and $\eta$ is the mass loading factor of galactic outflows. Provided the gas inflow rate into galaxies scales proportional to $M_{\rm halo}/t_{\rm ff}$, we recover Eq. (4) with $\epsilon\propto{}1/(1-R+\eta{})\sim{}1/\eta$ for low mass galaxies. The alternative choice $\dot{M}_{\rm gas,in}\propto{}\dot{M}_{\rm halo}$ is discussed further below.
@@ 3.3 The stellar mass – halo mass relation
One interesting aspect is the scaling of this predicted stellar fraction with halo mass at the massive end. Provided $M_{\rm star,min}\ll{}M_{\rm star}$, the halo mass scaling of the SFE is closely related to the scaling of the SHMR. Specifically, for a power-law ${\rm SFE}\propto{}M_{\rm halo}^{\alpha}$ (with $\alpha>-1$), $M_{\rm star}\propto{}M_{\rm halo}^{1+\alpha}$ and the stellar fraction $f_{\rm star}=M_{\rm star}/M_{\rm halo}\propto{}M_{\rm halo}^{\alpha}$, i.e., the ‘integrated SFE’ $f_{\rm star}$ has the same scaling with halo mass as the instantaneous SFE. A broken power-law fit to the $f_{\rm star}$ – halo mass relation obtained via Eq. (8) returns a high mass slope of $0.33$, in line with the expectation from the halo mass scaling of the SFE in FIREboxHR, see Table 1. This slope is, however, lower than the high mass slope ($\sim{}0.50$) of the actual $f_{\rm star}$ – halo mass relation in FIREboxHR.
@@ 3.4 A theoretical model for the UV luminosity function and density at $z>6$
As discussed in Section 3.1, FIREboxHR predicts UV luminosity densities at $z>10$ in good agreement with observational data by JWST. In this section, we introduce a simple model that provides a theoretical explanation for this result. Our approach builds on the ideas of previous work (Trenti et al., 2010; Tacchella et al., 2013; Mason et al., 2015; Furlanetto et al., 2017; Tacchella et al., 2018; Sabti et al., 2022; Harikane et al., 2022; Lovell et al., 2023; Ferrara et al., 2023) but differs in its use of a SFE – halo mass relation derived from a cosmological, hydrodynamical simulation. Given that the FIRE-2 physics model is not explicitly tuned to specific observables, this model is thus theoretical, and not empirical, in nature. This theoretical framework demonstrates that the gradual evolution of the UV luminosity density can be directly attributed to the baryonic processes, including stellar feedback and gas accretion, that take place in galaxies at high redshifts.
@@
It follows from Eq. (9) that $\mathcal{L}$ is also a log-normally distributed random variable with $\langle{}\mathcal{L}\rangle{}=\langle{}\mathcal{S}\rangle{}\dot{M}/\kappa$. In contrast, the rest-frame UV magnitude defined as
@@
We similarly obtain the number density of galaxies brighter than some luminosity $l$ or magnitude $m$ as
[…]
@@
where $p$ is the pdf of $\mathcal{M}$ with parameters given by (12). We evaluate the integrals in Eqs. (13, 15, 16) numerically with halo mass functions (HMFs) provided by the hmf module for Python (Murray et al., 2013; Murray, 2014). Appendix D provides an analytical solution of Eq. (16) in the case of a power-law HMF and SFE – halo mass relation.
The discussion so far has been quite general. The non-evolving SFE model makes the additional assumption, based on the results of the previous sections, that the average SFE and its scatter in halos of a given mass do not vary with redshift at $z>6$.
Figure 6: Evolution of the ultraviolet (UV) luminosity density down to a limiting observed rest-frame magnitude of -17 as predicted by FIREboxHR (blue solid line and error bars) and by a theoretical model with a non-evolving star formation efficiency (SFE) – halo mass relation inferred from the simulation (red line and shaded areas). The red line shows the UV luminosity density in galaxies brighter than -17 as predicted by the theoretical model for the same halo mass distribution as in FIREboxHR, while dark and light shaded areas indicate how the scatter in the SFE affects the UV luminosity in the given volume at the 68 and 95 percent confidence level. Uncertainties (68% confidence level) of the FIREboxHR predictions are calculated via bootstrapping and account for both the scatter in SFE and variations in halo numbers. Long and short dashed lines reproduce the predictions of empirical models by Tacchella et al. (2018) and Harikane et al. (2022), and large symbols with error bars show observational estimates by Bouwens et al. (2022), Donnan et al. (2023a); Donnan et al. (2023b), Harikane et al. (2023), Adams et al. (2024), Finkelstein et al. (2024) and Donnan et al. (2024). The theoretical model reproduces the evolution of the UV luminosity density found by FIREboxHR within the uncertainties.
@@
Figure 7: Predictions of the theoretical model for the cosmic ultraviolet (UV) luminosity density (left panel) and the corresponding galaxy number density (right panel) at $z\sim{}6-20$. The predictions (shown by solid lines) assume a non-evolving star formation efficiency – halo mass relation with mass-dependent scatter with parameters provided in Tables 1 and 2, a Tinker et al. (2008) halo mass function with the Behroozi et al. (2013) extension to high redshift, and Planck 2018 cosmology (Aghanim et al. 2020). Line colors indicate the adopted limiting magnitude (see legend). Symbols show observational estimates of the UV luminosity density and their uncertainties for $M_{\rm UV,lim}=-17$ (left panel) and galaxy number densities obtained by integrating the reported UV luminosity functions to $M_{\rm UV,lim}=-17$ (Bouwens et al., 2022; Donnan et al., 2023a; Donnan et al., 2023b; Harikane et al., 2023; Adams et al., 2024; Finkelstein et al., 2024; Donnan et al., 2024). The dotted line in the left panel is the UV luminosity density needed for a 21 cm absorption signal in the ‘minimal coupling’ regime (Madau, 2018). The UV luminosity density and number density predicted by the theoretical model for $M_{\rm UV,lim}=-17$ matches current observations and shows a steeper evolution at the highest redshift.
The theoretical model introduced in the previous section allows us to study the evolution of the UV luminosity density and UV LF during reionization and late Cosmic Dawn without being limited by simulation volume. To integrate eqs. (13, 16), we use the halo number densities provided by hmf version 3.4.4 for a Planck 2018 cosmology (Aghanim et al., 2020) with $h=0.6766$, $\Omega_{0}=0.3111$, $\Omega_{\rm b}=0.04897$, and $\sigma_{8}=0.8102$, a CAMB transfer function with [^8] Setting $k_{\rm max}$ to a sufficiently large value is critical as hmf’s default setting, at least in the code version we used, overestimates (e.g., by about 0.5 dex for $10^{8}\,M_{\odot}$ halos at $z=10$) the number density of low mass halos at high redshift. $k_{\rm max}=150$, and the Tinker et al. (2008) halo mass function with the Behroozi et al. (2013) extension to high redshift. To approximately account for the impact of baryonic effects on the halo mass (e.g., Sawala et al. 2013; Velliscig et al. 2014; Cui et al. 2014; Schaller et al. 2015; Beltz-Mohrmann & Berlind 2021), we reduce the halo masses provided by hmf by 0.07 dex, i.e., $M_{\rm halo}\rightarrow{}M_{\rm halo}^{\prime}=10^{-0.07}M_{\rm halo}$ and $n(>M_{\rm halo})\rightarrow{}n^{\prime}(>M_{\rm halo}^{\prime})=n(>M_{\rm halo})$. For the fiducial set up, we adopt the non-evolving SFE – halo mass relation with a mass-dependent scatter and with parameters provided in Tables 1 and 2 for dust-attenuated UV luminosities and halo mass binning. Variations of this set-up will be discussed in the text when needed. We stress that some model predictions involve extrapolations of the non-evolving SFE – halo mass ansatz to lower halo masses and higher redshifts than reliably probed by FIREboxHR. Additionally, the predictions assume no fundamental changes in baryonic physics and star formation processes at high $z$, such as contributions from Population III stars (see Section 3.7).
The left panel of Fig. 7 presents the UV luminosity density and UV LFs predicted by the theoretical model for $z\sim{}6-20$. The model accurately captures the observed evolution of the UV luminosity density over $z\sim{}6-13.5$. Notably, $\log{}\rho_{\rm UV}$ does not evolve linearly with redshift at $z\geq{}6$ (cf. Ceverino et al. 2024) showing instead a steeper decline in UV luminosity density towards higher redshifts. For instance, the UV luminosity density at $z=16$ (for a limiting magnitude of -17) is predicted to be about two orders of magnitude lower than at $z=11$, highlighting a challenge for observational studies targeting galaxies at those redshifts (e.g., Robertson et al. 2024). Extending this model speculatively to $z=20$ suggests a further reduction in luminosity density by an additional two orders of magnitude. Recent tentative observational evidence may suggest a sudden steepening of the UV luminosity density evolution between $z\sim{}12.5$ and $14.5$ (Robertson et al., 2024; Donnan et al., 2024). While a steeper evolution at higher $z$ is in line with our findings, our model predicts a gradual steepening over $z\sim{}6-15$, not a sudden transition between two separate evolutionary regimes.
Refer to caption Refer to caption
@@
Figure 8: Predictions of the theoretical model for the ultraviolet luminosity functions (UV LFs) at $z\sim{}6-20$. All predictions assume a non-evolving SFE – halo mass relation, a Tinker et al. (2008) halo mass function with the Behroozi extension to high redshift (Behroozi et al., 2013), and Planck 2018 cosmology (Aghanim et al. 2020). Solid lines of various colors indicate the model predictions for different redshifts. Dot-dashed and double dot-dashed lines show observational estimates of the UV LF at $z\sim{}6-12$ (Bouwens et al., 2022; Harikane et al., 2022; Donnan et al., 2023a; Harikane et al., 2023). (Left) Predictions for the fiducial model with a mass dependent scatter, see Table 2. Dotted lines (dashed lines) show fits of the model UV LFs to a modified Schechter (a double power law) function (Bouwens et al., 2017; Harikane et al., 2022). While both fitting functions describe the UV LFs of the theoretical model reasonably well, a double power law provides a significantly better fit, especially at $z\leq{}14$. (Right) Similar to left panel but for a mass-independent scatter of the SFE – halo mass relation. Solid (dotted, dashed) lines correspond to a scatter of 0.4 (0.2, 0.6) dex. A larger scatter increases the number density of UV luminous galaxies. In addition to a rapid decline in overall normalization, the theoretical model predicts a steepening in the faint-end of the UV LFs towards higher redshift. Fitting parameters for a modified Schechter function and double power law are provided in appendix E.
@@
The right panel of Fig. 7 show the number density of galaxies with rest-frame UV magnitudes brighter than a limiting value, $M_{\rm UV,lim}$, as predicted by our theoretical model. The predicted number densities match well observational estimates derived by integrating UV luminosity functions down to $M_{\rm UV,lim}=-17$. Deeper observations would dramatically increase the number density of detectable galaxies, e.g., by one order of magnitude when going from $M_{\rm UV,lim}=-17$ to $M_{\rm UV,lim}=-14$ at $z=6$. In contrast, the UV luminosity density is much less sensitive to the UV magnitude limit, increasing by only 0.3 dex at $z=6$ for the same change in $M_{\rm UV,lim}$.
The UV LFs predicted by the theoretical model are also in approximate agreement with observations, see Fig. 8. Deviations at the luminous end partly arise due to the lack of massive halos in FIREboxHR (given its modest box size) needed to constrain the high mass behavior of the SFE – halo mass relation. In addition, it is possible that FIREboxHR underestimates the scatter in the SFE in more massive halos. A larger scatter would indeed increase the number density of UV luminous galaxies as shown by the figure (see also Sun et al. 2023c). For instance, a SFE scatter of 0.6 dex (compared with the lower $\sim{}0.3$ dex found in FIREboxHR for $M_{\rm halo}\sim{}10^{10}-10^{11}\,M_{\odot}$ halos) would be sufficient to match the reported UV LF at $z=12$ by Harikane et al. (2023). However, applying the same scatter to $z=10$ would then lead to an overprediction of the observed UV LFs at this earlier redshift (Donnan et al., 2023a). The dependence of the UV LFs on the scatter is further discussed in Appendix D. A peculiar feature of the model with mass-dependent scatter is that it introduces a much more pronounced change in slope of the UV LF at moderately faint luminosities (e.g., near $-16.5$ at $z=6$) compared to the model with mass-independent scatter.
We fit the predicted UV LFs both with a modified Schechter function, see Section 4.3 in Bouwens et al. (2017), and a double power law (e.g., Harikane et al. 2022). The modified Schechter function includes a multiplicative term $10^{-0.4\delta{}(M_{\rm UV}+16)^{2}}$ to model a roll-over of the LF at the faint end. The fitting parameters are provided in appendix E. While not expected a priori, our model predictions are well fit by either functional form. In either case, and independently of the chosen scatter of the SFE – halo mass relation, fits to the UV LFs confirm the visually apparent decrease in the overall normalization and the steepening in the faint end slope with increasing redshift.
@@
The faint end slope of the UV LF, $\alpha$, is directly related to a combination of the low mass slope $\alpha_{1}$ of the SFE – halo mass relation, the slope $\alpha_{M}$ of the halo mass function over the relevant halo mass range, and the slope $\gamma$ of the halo growth rate. Specifically,
$\alpha=\frac{1+\alpha_{M}}{\alpha_{1}+\gamma}-1,$ (17)
assuming that the SFE, halo mass function, and halo growth rate follow power law relations with halo mass, and the scatter of the SFE is mass-independent, see Appendix D. For a given $\Lambda{}$CDM cosmology, the faint end slope of the UV LF thus directly probes the low mass slope of the SFE – halo mass relation. Alternatively, if the SFE – halo mass relation can be determined independently, then the faint end slope can serve as a constraint on the low-mass end of the halo mass function. In particular, if the SFE – halo mass relation is redshift independent, as suggested by FIREboxHR, the observed redshift evolution of $\alpha$ (e.g., Bouwens et al. 2012; Finkelstein et al. 2015; Bouwens et al. 2022) can be explained by the steepening (i.e., the decrease) of the slope $\alpha_{M}$ of the halo mass function for halos corresponding to moderately faint galaxies ($M_{\rm UV}\sim{}-18$ to $-16$) with increasing redshift.
@@
Finally, the figure analyzes how scatter in the SFE – halo mass relation, and the corresponding variability in the UV magnitudes, affects the cosmic UV luminosity density and the UV LFs. Specifically, we vary the mass-dependent scatter by ±$\pm{}0.2$ dex finding that the cosmic UV luminosity density is largely insensitive to this scatter. This finding is in qualitative agreement with a recent study by Gelli et al. (2024) who concluded that scatter alone does not explain the high UV luminosity density at $z>12$. Scatter in the SFE – halo mass relation has a significant impact, however, on the bright end of the UV LFs. As discussed before, a higher UV variability boosts the number density of UV bright galaxies given their much lower number density compared to faint ones. However, unless the scatter is increased to a value much larger than found in FIREboxHR, the number density of galaxies which dominate the cosmic UV luminosity density is only weakly affected. For a further discussion of the importance of the scatter on the UV LF, see Appendix D.
Our finding contrasts with alternative models that aim to explain the relatively shallow evolution of the UV luminosity density at high $z$ with a large and redshift-dependent scatter of the SFE (e.g., Shen et al. 2023b). We parametrize the SFE – halo mass relation in terms of both the mean relation and its log-normal scatter. As discussed in Section 3.4, an alternative apporach is to parametrize the SFE – halo mass relation in terms of the median relation and its scatter. However, for a log-normally distributed SFE, increasing the scatter while holding the median SFE fixed results in an increase in the mean SFE. A scatter that grows with redshift, while maintaining a fixed median SFE – halo mass relation, consequently raises the normalization of the average SFE – halo mass relation and causes a weaker decline in UV luminosity density with redshift. In contrast, our theoretical model successfully matches the observed evolution of the UV luminosity density without requiring an explicit redshift dependence of the average SFE – halo mass relation.